A rectangular lawn is measured at y meters wide by (2y+5) meters long. A) write an expression for the perimeter of the lawn. B) the perimeter is measured at 52 meters. Use this information to form an equation. C) solve this equation to find the length and width of the lawn
Answer by boilpoil(127)

A) Let X represent the perimeter of the lawn
X = 2(2y+5)+2y
= 4y+10+2y
= 6y+10

B) 6y+10=x and x=52
6y+10=52

C) 6y=42
y=7(meters)

Length of the lawn
=2[2(7)+5]
=2[14+5]
=2[19]
=38(meters)